OpenStudy (anonymous):

gcf 2^5x5x11 and 2^3x5^2x7 x is for times

OpenStudy (kc_kennylau):

\[2^5\times5\times11=\color{red}2\times\color{red}2\times\color{red}2\times2\times2\times\color{red}5\times11\]\[2^3\times5^2\times7=\color{red}2\times\color{red}2\times\color{red}2\times5\times\color{red}5\times7\]

OpenStudy (kc_kennylau):

can you see them? :)

OpenStudy (anonymous):

so it's 2^6 and 5^2?

OpenStudy (kc_kennylau):

wait, are you finding GCD or LCM?

OpenStudy (anonymous):

greatest common factor

OpenStudy (kc_kennylau):

can you see their common factors? how many 2 they have in common? how many 5 they have in common? how many 7 they have in common? how many 11 they have in common?

OpenStudy (anonymous):

three twos and the five

OpenStudy (kc_kennylau):

yep :D

OpenStudy (anonymous):

so 2^3*5? the * is for times

OpenStudy (anonymous):

woo!

OpenStudy (kc_kennylau):

exactly :D

OpenStudy (kc_kennylau):

for GCD, just look at their common factors :D

OpenStudy (anonymous):

so what about like -50m^4n^7 and 40m^2n^10

OpenStudy (kc_kennylau):

Break 'em all down to their components :DDDDD

OpenStudy (anonymous):

would that be -10m^2n^7?

OpenStudy (kc_kennylau):

exactly :D